# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ 3.2 Prior distribution
for $S_{t}\subseteq[p],$ $\mathbf{s}_{t}\in\prod_{i\in S_{t}}\mathcal{A}_{i}$ and $\beta_{t}\in\mathbb{R}.$
The parameters that need to be inferred are $T,$ $(S_{1},\mathbf{s}_{1},\beta_{1}),\ldots,(S_{T},\mathbf{s}_{T},\beta_{T})$ and $\sigma^{2}.$ First, we assume that prior of $(S_{t},\mathbf{s}_{t},\beta_{t}),t=1,\ldots,T$ are independent and identically distributed. In the following, we specify the prior distribution for $(S_{t},\mathbf{s}_{t},\beta_{t}),t=1,\ldots,T$ as well as $T$ and $\sigma^{2}$. Prior for $S$. We let $S$ follow the mixture distribution $\sum_{d=1}^{p}\omega_{d}\cdot\mbox{Uniform}\left\{{\rm power}([p],d)\right\}$ for $\omega_{d}\geq 0$ and $\sum_{d=1}^{p}\omega_{d}=1,$ where ${\rm power}([p],d)$ is the collection of all subsets of $[p]$ whose cardinality is $d.$ The weights $\omega_{d}$ are defined recursively as in BART , as follows. For a given positive integer $r,$ we let $p_{\text{split}}(d)=\alpha_{\text{split}}(1+d)^{-\gamma_{\text{split}}}$ for $\alpha_{\text{split}}\in(0,1)$ and $\gamma_{\text{split}}>0.$ Then, we set
@@
Basically, $\omega_{d}$ is decreasing in $d.$ The hyperparameters $\alpha_{\text{split}}$ and $\gamma_{\text{split}}$ control how fast $\omega_{d}$ decreases with $d$. That is, the probability $\text{Pr}(|S|>d)$ increases as $\alpha_{\text{split}}$ increases but decreases as $\gamma_{\text{split}}$ increases. Prior for $\mathbf{s}|S$. Conditional on $S$, $s_{j},j\in S$ are independent and uniformly distributed on $\mathcal{A}_{j},j\in S.$ In this paper, we set $\mathcal{A}_{j}=\{(x_{(i),j}+x_{(i-1),j})/2:i=2,\ldots,n\},$ where $x_{(i),j}$s are the order statistics of given data $\{x_{i,j},i=1,\ldots,n\}.$ Prior for $\beta$. We use a diffuse Gaussian prior
@@
for $\sigma_{\beta}^{2}>0.$ Prior for $\sigma^{2}$. As usual, we set
@@
for $v>0$ and $\lambda>0,$ where $IG(a,b)$ is the inverse gamma distribution with the shape parameter $a$ and scale parameter $b.$ Prior for $T$. We use the following distribution for the prior of $T:$
where $C_{*}>0$ and $T_{\max}\in\mathbb{N}_{+}$ are hyperparameters.
@@ 4 Posterior Sampling
In this section, we develop an MCMC algorithm for posterior sampling of ANOVA-BART. We use a Gibbs sampling algorithm to generate the parameters $T,(S_{1},\mathbf{s}_{1},\beta_{1}),\ldots,(S_{T},\mathbf{s}_{T},\beta_{T})$ and $\sigma^{2}$ from their conditional posterior distributions. For generating $(S_{t},\mathbf{s}_{t},\beta_{t})|\text{others},$ we modify the Metropolis-Hastings (MH) algorithm of BART (BART ). We can generate $\sigma^{2}|\text{others}$ easily since it is an inverse gamma distribution.
The hardest part is to generate $T|\text{others}.$ In fact, others in $T|\text{others}$ is even not well defined since $(S_{1},\mathbf{s}_{1},\beta_{1}),\ldots,(S_{T},\mathbf{s}_{T},\beta_{T})$ depend on $T.$ To resolve this problem and to make the convergence of the Gibbs sampler algorithm be faster, we rewrite $f$ in (8) by introducing the auxiliary binary randnom vector $\mathbf{z}=(z_{1},\ldots,z_{T_{\max}})^{\top}$ as
$f(\mathbf{x})=\sum_{t=1}^{T_{\max}}z_{t}\mathbb{T}(\mathbf{x}:S_{t},\mathbf{s}_{t},\beta_{t}).$ (10)
The prior distribution of $(z_{1},\ldots,z_{T_{\max}})|T$ is the uniform distribution on $\{0,1\}^{T_{\max}}$ subject to $\sum_{t=1}^{T_{\max}}z_{t}=T$ (i.e. the hyper-geometric distribution). We develop a Gibbs sampling algorithm based on the model (10).
@@
Let $\mathcal{D}=\{(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{n},y_{n})\}$ be given data which consist of $n$ pairs of observed covariate vector and response variable. For the likelihood, we assume that $y_{i}$s are realizations of $Y_{i}=f(\mathbb{x}_{i})+\epsilon_{i}$ conditional on $f$ and $\sigma^{2},$ where $\epsilon_{i}$ are independent Gaussian random noises with mean 0 and variance $\sigma^{2}.$
4.1 Sampling $(S_{t},\mathbf{s}_{t},\beta_{t})|\text{others}$
@@
If $z_{t}=0,$ the conditional posterior distribution is equal to the prior distribution of $(S_{t},\mathbf{s}_{t},\beta_{t})$. If $z_{t}=1$, the conditional posterior distribution depends on solely through $\text{Resid}_{t}$, where
$\text{Resid}_{t}=\left\{y_{i}-\sum_{k\neq t}z_{k}\mathbb{T}(\mathbb{x}_{i}:S_{k},\mathbf{s}_{k},\beta_{k}),i=1,\ldots,n\right\}$
@@
is the set of partial residuals derived from the current model that does not include the $t$th decision tree. Therefore, the sampling $(S_{t},\mathbf{s}_{t},\beta_{t})$ from the conditional posterior distribution is equal to sampling from $(S_{t},\mathbf{s}_{t},\beta_{t})|\text{Resid}_{t},\sigma^{2}.$ In turn, we decompose
$\pi\{S_{t},\mathbf{s}_{t},\beta_{t}|\text{Resid}_{t},\sigma^{2}\}=\pi\{S_{t},\mathbf{s}_{t}|\text{Resid}_{t},\sigma^{2}\}\pi\{\beta_{t}|\text{Resid}_{t},\sigma^{2},S_{t},\mathbf{s}_{t}\},$
and we first generate $(S_{t},\mathbf{s}_{t})$ and then generate $\beta_{t}$ from the corresponding conditional posterior distributions.
@@
For generating $(S_{t},\mathbf{s}_{t}),$ we use a modification of the MH algorithm used in BART (BART ). We use a proposal distribution similar to one used in bartmachine . A key difference is that we consider only identifiable binary-product trees defined in Section 3.1. As a proposal distribution in the MH algorithm, we consider the following three possible alterations of $(S_{t},\mathbf{s}_{t})$:
@@
  • GROW : adding an element $j^{\text{new}}$ selected randomly from $S_{t}^{c}$ into $S_{t}$ and choosing a split value for the newly selected element $j^{\text{new}}$ by selecting it randomly from $\mathcal{A}_{j^{\text{new}}}.$
@@
  • CHANGE : changing an element in $S_{t}$ by one selected from $S_{t}^{c}$ randomly and changing the split value accordingly.
The MH algorithm proposes $(S_{t}^{\text{new}},\mathbf{s}_{t}^{\text{new}})$ using one of GROW, PRUNE, and CHANGE with probability 0.28, 0.28 and 0.44, respectively (bartmachine ), and then accepts/rejects $(S_{t}^{\text{new}},\mathbf{s}_{t}^{\text{new}})$ according to the acceptance probability given in Section B.1 of Supplementary Material.
Generating $\beta_{t}$ from $\pi\{\beta_{t}|\text{Resid}_{t},\sigma^{2},S_{t},\mathbf{s}_{t}\}$ can be done easily since the conditional posterior distribution is again Gaussian. The details of $\pi\{\beta_{t}|\text{Resid}_{t},\sigma^{2},S_{t},\mathbf{s}_{t}\}$ are given in Section B.2 of Supplementary Material.
4.2 Sampling $\sigma^{2}|\text{others}$ and $\mathbf{z}|\text{others}$
Generating $\sigma^{2}$ can be done easily since the conditional posterior distribution of $\sigma^{2}$ is an inverse Gamma distribution whose details are given in Section B.3 of Supplementary Material.
For generating $\mathbf{z}|{\rm others},$ we use a MH algorithm. For a proposal $\mathbf{z}^{\text{new}},$ we choose $k$ randomly from $T_{\max}]$ and then let $z_{k}^{\text{new}}=1-z_{k}$ and $z_{j}^{\text{new}}=z_{j}$ for $j\neq k.$ Then, we accept $\mathbf{z}^{\text{new}}$ with the acceptance probability given in Section B.4 of Supplementary Material.
@@ 4.3 MCMC algorithm for ANOVA-BART
Algorithm 1 summarizes the proposed MCMC algorithm for ANOVA-BART. The MCMC algorithm is similar to that for BART(BART ), but it has the additional sampling of $\mathbb{z}$. The sampling of $\mathbb{z}$ is not needed for BART since the number of decision trees $T$ is fixed.
@@
Input: $T_{\max}$ : maximum number of decision trees, $M$ : number of MCMC iteration
1:for i : 1 to M do
2:  for t : 1 to $T_{\text{max}}$ do
3:   if $z_{t}=1$ then
4:    $(S_{t},\mathbf{s}_{t})\sim\text{MH}_{S_{t},\mathbf{s}_{t}}(\text{Resid}_{t},\sigma^{2})$
5:    $\beta_{t}\sim\pi\{\beta_{t}|\text{Resid}_{t},\sigma^{2},S_{t},\mathbf{s}_{t}\}$
6:   else
7:    $(S_{t},\mathbf{s}_{t},\beta_{t})\sim\pi\{S_{t},\mathbf{s}_{t},\beta_{t}\}$
8:   end if
@@
10:  $\mathbf{z}\sim\text{MH}_{\mathbf{z}}(\mathcal{D},\{(S_{t},\mathbf{s}_{t})\}_{t=1}^{T_{\max}},\sigma^{2})$
11:  $\sigma^{2}\sim\pi\{\sigma^{2}|\text{others}\}$
12:  return  $\{\mathbb{T}(\cdot:S_{t},\mathbf{s}_{t},\beta_{t})\}_{t:z_{t}=1},\sigma^{2}$
13:end for
Remark 4.1.
In the algorithm, we do not need to generate $(S_{t},\mathbf{s}_{t},\beta_{t})$ for $z_{t}=0$ at every iteration. Instead, we generate $(S_{t},\mathbf{s}_{t},\beta_{t})$ only when $z_{t}=0$ but $z_{t}^{\text{new}}=1$ to calculate $\pi\{\mathbf{z}^{\text{new}}|\text{others}\}$
@@ 5 Posterior concentration rate
In this section, we study theoretical properties of ANOVA-BART. In particular, we drive the posterior concentration rate of ANOVA-BART when the true model is smooth (i.e. Hölder smooth). Our posterior concentration rate is the same as the minimax optimal rate (up to a logarithm term) and adaptive to the smoothness of the true function. Moreover, we derive the posterior concentration rate of each interaction, which could be utilized to screen out unnecessary interactions out after posterior computation.
We consider the random-$X$ design. That is, the input-output pairs $(\mathbb{x}_{1},y_{1}),\ldots,(\mathbb{x}_{n},y_{n})$ are assumed to be a realization of $(\mathbf{X}_{1},Y_{1}),\ldots,(\mathbf{X}_{n},Y_{n})$ that are independent copies of $(\mathbb{X},Y)$ whose distribution $\mathbb{P}_{0}$ is given as
$\mathbb{X}\sim\mathbb{P}_{\mathbb{X}}\quad\text{and}\quad Y|\mathbb{X}=\mathbb{x}\sim N(f_{0}(\mathbb{x}),\sigma_{0}^{2}).$
Here, $f_{0}$ and $\sigma_{0}^{2}$ are the true regression function and the true variance, respectively. Let $\mathbf{X}^{(n)}=(\mathbf{X}_{1},...,\mathbf{X}_{n})^{\top}$ and $Y^{(n)}=(Y_{1},...,Y_{n})^{\top}$. The main reason of considering the random-$X$ design rather than the fixed-$X$ design is to apply the populational identifiability condition ($\mathbb{P}_{\mathbf{X}}$-identifiability condition) to the functional ANOVA decomposition of $f_{0}.$ Discussions about the fixed-$X$ design are given in Section G of Supplementary Material.
To derive the posterior concentration rate of ANOVA-BART, as usual, we will check the sufficient conditions given by ghosal2000convergence . The most technically difficult part is to study the approximation property of the sum of identifiable binary-product trees with bounded heights. Theorem C.1 in Section C.3 of Supplementary Material is the main approximation theorem for this purpose.
For technical reasons, we consider the truncated prior $\pi_{\xi}$ for a large positive constant $\xi$ defined as $\pi_{\xi}\{\cdot\}\propto\pi\{\cdot\}\mathbb{I}(\|f\|_{\infty}\leq\xi,1/\xi\leq\sigma^{2}\leq\xi),$ where $\pi\{\cdot\}$ is the prior introduced in Section 3.2. For posterior sampling with this prior, we generate samples of $(f(\cdot),\sigma^{2})$ by the MCMC algorithm developed in Section 4 and only accept samples satisfying $\|f\|_{\infty}\leq\xi$ and $1/\xi\leq\sigma^{2}\leq\xi.$ We denote $\pi_{\xi}\{\cdot|\mathcal{D}\}$ the corresponding posterior.

### v3

@@ 3.2 Prior distribution
for $S_{t}\subseteq[p],$ $\mathbf{s}_{t}\in\prod_{i\in S_{t}}\mathcal{A}_{i}$ and $\beta_{t}\in\mathbb{R}.$ The parameters to be inferred are $T$ as well as $(S_{1},\mathbf{s}_{1},\beta_{1}),\ldots,(S_{T},\mathbf{s}_{T},\beta_{T})$ and $\psi.$ In the following, we specify the prior distribution $T,$ for $(S_{t},\mathbf{s}_{t},\beta_{t}),t=1,\ldots,T$ and $\psi$. First, we assume that a priori $(S_{t},\mathbf{s}_{t},\beta_{t}),t=1,2,\ldots$ are independent and identically distributed. Prior for $S$. We let $S$ follow the mixture distribution $\sum_{d=1}^{p}\omega_{d}\cdot\mbox{Uniform}\left\{{\rm power}([p],d)\right\}$ for $\omega_{d}\geq 0$ and $\sum_{d=1}^{p}\omega_{d}=1$. The weights $\omega_{d}$ are defined recursively as in Chipman et al. (2010), as follows. For a given positive integer $d,$ we let $p_{\text{split}}(d)=\alpha_{\text{split}}(1+d)^{-\gamma_{\text{split}}}$ for $\alpha_{\text{split}}\in(0,1)$ and $\gamma_{\text{split}}>0.$ Then, we set
@@
Basically, $\omega_{d}$ is decreasing in $d.$ The hyperparameters $\alpha_{\text{split}}$ and $\gamma_{\text{split}}$ control how fast $\omega_{d}$ decreases with $d$. That is, the probability $\text{Pr}(|S|>d)$ increases as $\alpha_{\text{split}}$ increases but decreases as $\gamma_{\text{split}}$ increases. Prior for $\mathbf{s}|S$. Conditional on $S$, $\{s_{j},j\in S\}$ are independent and uniformly distributed on $\mathcal{A}_{j},j\in S.$ In this paper, we set
$\mathcal{A}_{j}=\{(x_{(i),j}+x_{(i-1),j})/2:i=2,\ldots,n\},$ (7)
where $x_{(i),j}$s are the order statistics of given data $\{x_{i,j},i=1,\ldots,n\}.$ Prior for $\beta$. We use a diffuse Gaussian prior
@@
for $\sigma_{\beta}^{2}>0.$ Prior for $T$. We use the following distribution for the prior of $T:$
where $C_{*}>0$ and $T_{\max}\in\mathbb{N}_{+}$ are hyperparameters. Prior for the nuisance parameter $\psi$. For a model having the nuisance parameter, we only consider the Gaussian regression model: $p_{f(\mathbf{x}),\psi}(y)\propto\exp(-(y-f(\mathbf{x}))^{2}/2\sigma^{2}),$ where $\psi=\sigma^{2},$ and we let
@@
for $v>0$ and $\lambda>0$ a priori, where $IG(a,b)$ is the inverse gamma distribution with the shape parameter $a$ and scale parameter $b.$
@@ 4 Posterior Sampling
In this section, we develop an MCMC algorithm for posterior sampling of ANOVA-BART by modifying the MCMC algorithms of BART (Chipman et al. (2010), Kapelner and Bleich (2016), Linero and Yang (2018)). Note that, in contrast to Generalized BART ((Linero, 2025)), the MCMC algorithm for ANOVA-BART does not require reversible jump MCMC when updating the tree structure, since the dimension of the height parameter is fixed at one and remains unchanged across different tree structures.
Basically, we use a Gibbs sampling algorithm to generate the parameters $T,\{(S_{t},\mathbf{s}_{t},\beta_{t}),t\in[T]\}$ and nuisance parameter $\psi$ from their conditional posterior distributions. That is, our MCMC algorithm proceeds with the following three steps:
  1. 1.
    Update $T$.
  2. 2.
    Update $(S_{t},\mathbf{s}_{t},\beta_{t})$ for $t=1,\ldots,T.$
  3. 3.
    Update the nuisance parameter $\psi$.
Let $\Theta=(T,S_{1},\mathbf{s}_{1},\beta_{1},\ldots,S_{T},\mathbf{s}_{T},\beta_{T},\psi)$. Define $P_{\boldsymbol{\mathfrak{w}}}$ as a multinomial distribution with one trial on $[p]$ and probability vector $\boldsymbol{\mathfrak{w}}=(\mathfrak{w}_{1},\ldots,\mathfrak{w}_{p})^{\top}$, where $\mathfrak{w}_{j}\propto\sum_{t=1}^{T}\mathbb{I}(j\in S_{t})$ for $j\in[p]$.
4.1 Updating $T$
We employ the following proposal distribution in the MH algorithm. Given $T$, we propose $T^{\text{new}}=T-1$ with probability $T/T_{\max}$, and $T^{\text{new}}=T+1$ with probability $1-T/T_{\max}$.
Case of $T^{\text{new}}=T-1$.
We obtain $\Theta^{\text{new}}$ by deleting one element randomly from $\{(S_{t},\mathbf{s}_{t},\beta_{t}),t=1,\ldots,T\}.$
Case of $T^{\text{new}}=T+1$.
A central step is to construct a suitable proposal for the new tree, i.e, $(S_{T+1}^{\text{new}},\mathbf{s}_{T+1}^{\text{new}},\beta_{T+1}^{\text{new}})$. For this purpose, we first propose $S_{T+1}^{\text{new}}$, and then $\mathbf{s}_{T+1}^{\text{new}}$ and $\beta_{T+1}^{\text{new}}$ are drawn from their prior distributions given $S_{T+1}^{\text{new}}.$ We devise a proposal distribution of $S_{T+1}^{\text{new}}$ carefully to ensure that the MCMC algorithm can explore higher order interactions efficiently.
Let $M$ be a given positive integer. First of all, the proposal distribution for $S_{T+1}^{\text{new}}$ consists of the following two alternatives:
  • Random : With probability $M/(M+T),$ generate $S_{T+1}^{\text{new}}$ from the prior distribution,
  • Stepwise: With probability $T/(M+T),$ set $S_{T+1}^{\text{new}}=S_{t^{*}}\cup\{j_{t^{*}}\},$ where $t^{*}\sim\text{Uniform}[T]$ and $j_{t^{*}}\sim P_{\boldsymbol{\mathfrak{w}}}(\cdot|\cdot\in S_{t^{*}}^{c}).$
The stepwise move is designed to ensure that the order of a newly added tree is always one higher than that of an existing tree in the current ensemble, thereby encouraging the MCMC algorithm to explore higher-order interactions more frequently. In addition, we use $P_{\boldsymbol{\mathfrak{w}}}$ to increase the acceptance rate.
4.2 Updating $(S_{t},\mathbf{s}_{t},\beta_{t})$ for t=1,,Tt=1,...,T
@@ 4.1 Updating $T$
Let $\{(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{n},y_{n})\}$ be given data which consist of $n$ pairs of observed covariate vector and response variable. For the likelihood, we assume that $y_{i}$s are realizations of $Y_{i}|\mathbf{x}_{i}\sim\mathbb{P}_{f(\mathbf{x}_{i}),\psi}$ conditional on $f$ and $\psi$. Let $\mathbf{y}^{(n)}:=(y_{1},...,y_{n})$.
Updating $(S_{t},\mathbf{s}_{t})$.
@@
The conditional posterior distribution $\pi\{S_{t},\mathbf{s}_{t}|\text{others}\}$ depends on solely through $\lambda_{t},\mathbf{y}^{(n)},\beta_{t}$ and $\psi$, where
$\lambda_{t}=\left\{\sum_{k\neq t}\mathbb{T}(\mathbb{x}_{i}:S_{k},\mathbf{s}_{k},\beta_{k}),\>i=1,\ldots,n\right\}$
@@
is the set of partial predictive values derived from the current model that does not include the $t$th decision tree. Therefore, sampling $(S_{t},\mathbf{s}_{t})$ from the conditional posterior distribution is equivalent to sampling from the distribution $\pi\{S_{t},\mathbf{s}_{t}\mid\lambda_{t},\mathbf{y}^{(n)},\beta_{t},\psi\}$.
@@
For updating $(S_{t},\mathbf{s}_{t}),$ we use a modification of the MH algorithm used in BART (Chipman et al. (2010)). We use a proposal distribution similar to one used in Kapelner and Bleich (2016); Linero (2018a). A key difference is that we consider only identifiable binary-product trees defined in Section 3.1. As a proposal distribution in the MH algorithm, we consider the following three possible alterations of $(S_{t},\mathbf{s}_{t})$:
@@
  • GROW : adding an element $j^{\text{new}}$ selected randomly based on $P_{\boldsymbol{\mathfrak{w}}}(\cdot|\cdot\in S_{t}^{c})$ into $S_{t}$ and choosing a split value for the newly selected element $j^{\text{new}}$ by selecting it randomly from $\mathcal{A}_{j^{\text{new}}}.$
@@
  • CHANGE : changing an element in $S_{t}$ by one randomly selected from $S_{t}^{c}$ based on $P_{\boldsymbol{\mathfrak{w}}}(\cdot|\cdot\in S_{t}^{c})$ and changing the split value accordingly.
The MH algorithm proposes $(S_{t}^{\text{new}},\mathbf{s}_{t}^{\text{new}})$ using one of GROW, PRUNE, and CHANGE with probability 0.28, 0.28 and 0.44, respectively (Kapelner and Bleich (2016)), and then accepts/rejects $(S_{t}^{\text{new}},\mathbf{s}_{t}^{\text{new}})$ according to the acceptance probability given in Section A.2 of Supplementary Material.
Updating $\beta_{t}$.
We update $\beta_{t}$ via the MH algorithm with the Langevin dynamics (Rossky et al. (1978)). The details are given in Section A.3 of Supplementary Material.
4.3 Updating nuisance parameter $\psi$
Recall that $\psi=\sigma^{2},$ where $\sigma^{2}$ is the variance of the noise in the Gaussian regression model. Generating $\sigma^{2}$ can be done easily since the conditional posterior distribution of $\sigma^{2}$ is an inverse Gamma distribution whose details are given in Section A.4 of Supplementary Material.
@@ 4.4 MCMC algorithm for ANOVA-BART
Algorithm 1 summarizes the proposed MCMC algorithm for ANOVA-BART. Note that the sampling of $T$ is not needed for BART since the number of decision trees $T$ is fixed.
@@
Input: $T$ : initial number of decision trees, $\mathbb{M}$ : number of MCMC iterations
1:for i : 1 to $\mathbb{M}$ do
2:  $T\sim\text{MH}_{T}(\{S_{t},\mathbf{s}_{t},\beta_{t}\}_{t=1}^{T},\psi)$
3:  for t : 1 to $T$ do
4:   $(S_{t},\mathbf{s}_{t})\sim\text{MH}_{S_{t},\mathbf{s}_{t}}(\lambda_{t},\mathbf{y}^{(n)},\beta_{t},\psi)$
5:   $\beta_{t}\sim\text{MH}_{\beta_{t}}(\lambda_{t},\mathbf{y}^{(n)},S_{t},\psi)$
6:  end for
7:  $\psi\sim\pi\{\psi|\text{others}\}$
8:  return $\{\mathbb{T}(\cdot:S_{t},\mathbf{s}_{t},\beta_{t})\}_{t=1}^{T},\psi$
@@ 5 Posterior concentration rate
In this section, we study theoretical properties of ANOVA-BART. In particular, we drive the posterior concentration rate of ANOVA-BART when the true model is smooth (i.e., Hölder smooth). Our posterior concentration rate is the same as the minimax optimal rate (up to a logarithm term) and adaptive to the smoothness and the maximum order of signal interactions for the true function. Moreover, we derive the posterior concentration rate of each component, which can be utilized to screen out unnecessary components after posterior computation.
For technical simplicity, we consider the random-$X$ design and fix the nuisance parameter $\psi$ as 1. That is, the input-output pairs $(\mathbb{x}_{1},y_{1}),\ldots,(\mathbb{x}_{n},y_{n})$ are assumed to be a realization of $(\mathbf{X}_{1},Y_{1}),\ldots,(\mathbf{X}_{n},Y_{n})$ that are independent copies of $(\mathbb{X},Y)$ whose distribution $\mathbb{P}_{0}$ is given as
$\mathbb{X}\sim\mathbb{P}_{\mathbb{X}}\quad\text{and}\quad Y|\mathbb{X}=\mathbb{x}\sim\mathbb{P}_{f_{0}(\mathbb{x}),1},$
where $f_{0}$ is the true regression function. Let $\mathbf{X}^{(n)}=(\mathbf{X}_{1},...,\mathbf{X}_{n})^{\top}$ and $Y^{(n)}=(Y_{1},...,Y_{n})^{\top}$. The main reason of considering the random-$X$ design rather than the fixed-$X$ design is to apply the populational identifiability condition ($\mathbb{P}_{\mathbf{X}}$-identifiability condition) to the functional ANOVA decomposition of $f_{0}.$ Discussions about the fixed-$X$ design are given in Section G of Supplementary Material.
To derive the posterior concentration rate of ANOVA-BART, as usual, we will check the sufficient conditions given by Ghosal and van der Vaart (2007). The most technically difficult part is to study the approximation property of the sum of identifiable binary-product trees with bounded heights. Theorem B.1 in Section B.3 of Supplementary Material is the main approximation theorem for this purpose.
For technical reasons, we consider the truncated prior $\pi_{\xi}$ for a large positive constant $\xi$ defined as $\pi_{\xi}\{\cdot\}\propto\pi\{\cdot\}\mathbb{I}(\|f\|_{\infty}\leq\xi),$ where $\pi\{\cdot\}$ is the prior introduced in Section 3.2. For posterior sampling with this prior, we generate samples of $f(\cdot)$ by the MCMC algorithm developed in Section 4 and only accept samples satisfying $\|f\|_{\infty}\leq\xi$. We denote $\pi_{\xi}\{\cdot|\mathbf{X}^{(n)},Y^{(n)}\}$ the corresponding posterior.